OpenStudy (anonymous):

what is the mass in grams of 13.2 amu?

OpenStudy (rogue):

Well, there are 6.02 x 10^23 amu is 1 gram. So can you figure out how many grams are in 13.2 amu?

OpenStudy (anonymous):

1 amu = 1/12* mass of C-12atom=1/12*12/6.02*10^23= 1 amu =1.66*10-24 g

OpenStudy (rogue):

\[13.2 amu \times \frac {1 gram}{6.02 \times 10^{23} amu} = 2.192 \times 10^{-23} grams\]